# Multi-person Physics-based Pose Estimation for Combat Sports

### v1

@@ 2 Related Work
Multi-person 3D pose estimation in sports poses significant challenges due to fast and unpredictable movements, frequent occlusions, and the need for fine-grained motion capture under high-intensity conditions [33]; early methods relied on controlled environments with multi-camera setups or marker-based motion capture systems that, while accurate, proved costly and inflexible for real-world scenarios. Recent advances have shifted toward leveraging monocular or multi-view video feeds in unconstrained settings to enable applications such as player tracking, action recognition, and performance analysis in sports like basketball, soccer [25], and tennis [51, 6, 64]. To address issues arising from frequent occlusions and complex player interactions, many approaches now integrate spatio-temporal modules or graph-based methods to capture both short-term and long-term dynamics in high-speed sports sequences [55, 66], while Transformer-based architectures have shown promise in effectively modeling global context in dense player formations [16]. The emergence of large-scale sports datasets with rich annotations further facilitates improved generalization under domain-specific constraints [29, 6], spurring the development of specialized datasets and systems such as SportsPose [23], which features markerless motion capture of five high-speed sports validated against commercial marker-based systems, and AthletePose3D [61], which captures extreme dynamics across 12 sport motion types and approximately 1.3 million frames to achieve a reduction in pose error by nearly 69% (from 214 mm to 65 mm). For team sports, WorldPose [25] provides a large-scale multi-person dataset from the 2022 FIFA World Cup that captures 3D poses and trajectories of entire soccer teams using a static multi-camera setup, while tailored applications like Reconstructing NBA Players [69] demonstrate the integration of computer vision with domain-specific data to achieve high-resolution 3D mesh reconstruction from single broadcast images despite challenges such as complex poses, motion blur, and occlusions. In dense crowd scenarios, robust multi-view geometry frameworks that combine probabilistic reasoning and cross-view matching have proven effective [5]. In this work, we focus on using multi-person physics-based pose estimation of elite boxers captured during sparring sessions to robustly recover detailed 3D kinematics under challenging occlusion and rapid motion conditions. Our approach uniquely integrates physics simulation with advanced deep learning, enabling precise biomechanical analysis and offering new insights for performance optimization in competitive boxing.
@@ 3 Methodology
Figure 2 illustrates the overall architecture of our proposed framework, comprising four key stages: multi-view tracking, weighted triangulation, kinematic optimization, and dynamics-based trajectory refinement. Each stage is detailed in Sections 3.1 through 3.4.
[…]
@@ 3.1 Multi-frame Multi-view Tracking IDs
[…]
Figure 3: Illustration of the epipolar geometry distance for a pair of cross view 2D positions.
@@ 3.4 Multi-person Dynamic Optimization
Figure 4: Illustration of SMPL to physics-based humanoid conversion for varying body shapes.
@@ Conclusion.
Supplementary Material
5.1 Evaluation Metrics
We evaluate our approach using two main categories of metrics: pose metrics and physics metrics.
Pose Metrics. These measure pose accuracy and smoothness:
  • WA-MPJPE (World-Aligned MPJPE). In some works, this refers to an MPJPE calculation where the predicted and ground-truth poses are aligned in the global (or “world”) coordinate system by a single rigid transformation (often estimated via a Procrustes-like method) once for the entire sequence. It does not allow a different transformation per frame.
  • W-MPJPE (World PA First-MPJPE). This measures MPJPE after aligning only the first frame of the prediction and ground truth in the global coordinate system. The remaining frames are evaluated without further alignment, highlighting any drift that occurs over time.
  • PA-MPJPE (Procrustes Aligned MPJPE). Each frame of the predicted pose is rigidly aligned to the corresponding ground-truth pose on a per-frame basis before computing MPJPE. This alignment factors out any frame-by-frame rigid transformations, emphasizing errors in limb positions or shape.
  • Acceleration Error (Acc. Error). Evaluates discrepancies in acceleration between the predicted motion and ground truth, indicating whether the motion is jittery or overly smoothed.
Physics Metrics. These assess how physically plausible a motion sequence is:
  • Foot Skating (Skating). Evaluates the extent to which feet slide on the ground when they should remain still.
  • Ground Penetration (Gnd Pen.). Measures whether any body part intersects or penetrates the ground plane.
  • Inter-Person Penetration (Pen.). Assesses collisions among different individuals by computing signed distance function (SDF) values. For each person, the SDF values from penetrating vertices of others are accumulated over all frames and averaged per person. This metric is measured in millimeters, whereas the other physics metrics are measured in meters.
Below are further details on the individual metrics:
5.1.1 Percentage of Correct Parts (PCP)
The PCP metric quantifies the percentage of body parts whose endpoints are correctly localized. Let the ground-truth body part endpoints be $\mathbf{s}_{c}$ and $\mathbf{e}_{c}$, and let $\mathbf{s}_{n}$ and $\mathbf{e}_{n}$ be their estimated counterparts. The part is considered correct if:
$|\mathbf{s}-\mathbf{s}_{c}|+|\mathbf{e}-\mathbf{e}_{c}|\;\leq\;|\mathbf{s}_{n}-\mathbf{e}_{n}|\,.$ (8)
5.1.2 Mean Per Joint Position Error (MPJPE)
The MPJPE metric computes the mean Euclidean distance between predicted and ground-truth 3D joint positions:
$\mathrm{e}_{MPJPE}=\mathrm{mean}\bigl{\|}\hat{\mathbf{p}}_{t}-\mathbf{p}_{t}^{GT}\bigr{\|}.$ (9)
5.1.3 Foot Floating and Sliding
Foot Floating.
We compare the vertical position of the foot in the predicted motion to the ground truth:
$\mathrm{e}_{foot,z}=\mathrm{mean}\bigl{\|}\hat{\mathbf{p}}_{foot,z}-\mathbf{p}_{foot,z}^{GT}\bigr{\|}.$ (10)
Foot Sliding.
We compare the foot’s horizontal velocity (on the XY plane) in the predicted motion to the ground truth:
$\mathrm{e}_{foot,v_{xy}}=\mathrm{mean}\bigl{\|}\Delta\hat{\mathbf{p}}_{foot,{v_{xy}}_{t}}-\Delta\mathbf{p}_{foot,{v_{xy}}_{t}}^{GT}\bigr{\|}.$ (11)
5.1.4 Motion Smoothness
Smoothness is measured by (1) computing MPJPE on a 15-joint reduced skeleton plus the root, and (2) comparing keypoint velocities:
$\hat{\mathbf{Jit}}=\bigl{\|}\hat{\mathbf{p}}_{t}-\hat{\mathbf{p}}_{t-1}\bigr{\|},$ (12)
$\mathbf{Jit}^{GT}=\bigl{\|}\mathbf{p}^{GT}_{t}-\mathbf{p}^{GT}_{t-1}\bigr{\|},$ (13)
$\mathrm{e}_{smooth}=\mathrm{mean}\,\bigl{\|}\mathbf{Jit}^{GT}-\hat{\mathbf{Jit}}\bigr{\|}.$ (14)
5.2 Ablation Study
We conduct ablation experiments to illustrate the impact of key components in our framework.
Number of Views.
Reducing camera views typically degrades 3D reconstruction due to weaker triangulation. Nevertheless, our method retains robustness in Table 6, even when views are sparse.
Triangulation Filtering.
Spline smoothing and an Extended Kalman Filter (EKF) on keypoints significantly denoise the initial linear triangulation, improving subsequent kinematic optimization (Table 6).
Prior (GMM + Vposer).
A Gaussian Mixture Model (GMM) prior in combination with Vposer narrows the pose parameter search space to more plausible solutions.
Regularization.
Pose regularization encourages smooth transitions between frames, reducing sudden jumps or unnatural artifacts.
Table 6: Ablation Study on Shelf Dataset using PCP metric.
Method Actor 1 Actor 2 Actor 3 Average
Baseline 99.8 97.6 98.6 98.6
Using 3 Views 98.3 98.1 98.6 98.3
Using 2 Views 97.6 48.1 96.4 80.7
No Filtering for Triangulation 99.8 94.9 97.9 97.5
No Prior 99.8 96.2 98.1 98.0
No 2D loss 91.6 76.2 81.2 83.0
No 3D loss 99.8 96.2 98.1 98.0
No pose regularization 99.8 95.7 98.5 98.0
Table 7: Configuration of the Humanoid used in physics optimization.
Joint Force Range Angle Range (rad) Joint Type No. Axes Damping
Head -40 to 40 -0.524 to 0.524 Hinge 3 20
Clavicle -80 to 80 -0.349 to 0.349 Hinge 2 20
Femur -300 to 300 -2.792 to 1.221 Hinge 3 20
Tibia -160 to 160 0.01 to 2.967 Hinge 1 16
Foot -120 to 120 -0.785 to 1.222 Hinge 2 12
Hand -20 to 20 -1.57 to 1.57 Hinge 2 1
Humerus -120 to 120 -1.571 to 1.571 Hinge 3 6
Lower Back -300 to 300 -0.523 to 0.523 Hinge 3 24
Lower Neck -120 to 120 -0.5236 to 0.5236 Hinge 3 40
Radius -90 to 90 -2.967 to 0.174 Hinge 1 5
Toes -20 to 20 1.570 to 0.349 Hinge 1 1
Wrist -20 to 20 -3.14159 to 0 Hinge 1 1
Root - - Free - -
5.3 CMU Humanoid Details
Table 7 shows the joint configurations for the CMU humanoid employed in physics optimization. Although the physics engine treats joints with three rotation axes as ball joints, we constrain these axes to mimic the behavior of three consecutive hinge joints, providing finer rotational control.
In summary, our pipeline combines noise-robust triangulation, strong pose priors, and physics-based optimization to yield smooth, physically plausible human motions under diverse scenarios, including sparse camera setups and noisy estimates.

### v2

@@ 2 Related Work
Multi-person 3D pose estimation in sports poses significant challenges due to fast and unpredictable movements, frequent occlusions, and the need for fine-grained motion capture under high-intensity conditions [5]; early methods relied on controlled environments with multi-camera setups or marker-based motion capture systems that, while accurate, proved costly and inflexible for real-world scenarios. Recent advances have shifted toward leveraging monocular or multi-view video feeds in unconstrained settings to enable applications such as player tracking, action recognition, and performance analysis in sports like basketball, soccer [24], and tennis [49]. To address issues arising from frequent occlusions and complex player interactions, many approaches now integrate spatio-temporal modules or graph-based methods to capture both short-term and long-term dynamics in high-speed sports sequences [35, 56], while Transformer-based architectures have shown promise in effectively modeling global context in dense player formations [43]. The emergence of large-scale sports datasets with rich annotations further facilitates improved generalization under domain-specific constraints [22], spurring the development of specialized datasets and systems such as SportsPose [22], which features markerless motion capture of five high-speed sports validated against commercial marker-based systems, and AthletePose3D [58], which captures extreme dynamics across 12 sport motion types and approximately 1.3 million frames to achieve a reduction in pose error by nearly 69% (from 214 mm to 65 mm). For team sports, WorldPose [24] provides a large-scale multi-person dataset from the 2022 FIFA World Cup that captures 3D poses and trajectories of entire soccer teams using a static multi-camera setup, while tailored applications like Reconstructing NBA Players [64] demonstrate the integration of computer vision with domain-specific data to achieve high-resolution 3D mesh reconstruction from single broadcast images despite challenges such as complex poses, motion blur, and occlusions. In dense crowd scenarios, robust multi-view geometry frameworks that combine probabilistic reasoning and cross-view matching have proven effective [6]. In this work, we focus on using multi-person physics-based pose estimation of elite boxers captured during sparring sessions to robustly recover detailed 3D kinematics under challenging occlusion and rapid motion conditions. Our approach uniquely integrates physics simulation with advanced deep learning, enabling precise biomechanical analysis and offering new insights for performance optimization in competitive boxing.
@@ 3 Methodology
Figure LABEL:fig:3d illustrates the overall architecture of our proposed framework, comprising four key stages: multi-view tracking, weighted triangulation, kinematic optimization, and dynamics-based trajectory refinement. Each stage is detailed in Sections 3.1 through 3.4.
@@ 3.1 Multi-frame Multi-view Tracking IDs
Figure 3: Illustration of the epipolar geometry distance for a pair of cross view 2D positions
@@ 3.4 Multi-person Dynamic Optimization
Figure 4: Illustration of SMPL to physics-based humanoid conversion for varying body shapes